# Symbolic recursion method for strongly correlated fermions in two and three dimensions

### v1

@@ Abstract
In Fig. 3, we present the infinite-temperature diffusion constants for two- and three-dimensional $t$-$V$ and Hubbard models as functions of the interaction strength $V$ (calculations are detailed in [supplement]). Perturbation theory predicts $D\sim 1/V^{2}$ at small $V$ [Kiely_2021_Transport, Vucicevic_2023_Charge]. Surprisingly, we find that this scaling extends to quite large values of $V$, far beyond the expected range of applicability of the perturbation theory.
There exists a compelling experimental [Brown_2019_Bad] and theoretical [Brown_2019_Bad, Mousatov_2019_bad, Vranic_2020_Charge, Kiely_2021_Transport, Vucicevic_2023_Charge] evidence that the linear scaling of conductivity with $\beta$, as expressed in Eq. (10), remains valid down to relatively low temperatures on the order of 1 or even less. Combined with this observation, our calculations yield results for the dc conductivity across a wide range of interaction strengths and temperatures.
Our result for the diffusion constant in the 2D Hubbard model is consistent with a recent tour de force perturbative calculation reported in Ref. [Kovacevic_2025_Toward]. Notably, the latter study has culminated several earlier attempts [Kiely_2021_Transport, Vucicevic_2023_Charge] that had missed certain relevant leading-order perturbative contributions. The intricacy of this perturbative calculation underscores the nontrivial nature of the hight-temperature quantum transport.
Majorana Strings
Mention similarity to Pauli Strings basis, but with the detail that now we need to canonicalize the strings and that it does not increases the memory but increases time. Finally, we discuss technical aspects of calculating the moments (5). To compute the $n$-fold nested commutator $H(V),J]^{(n)}$, we employ an operator basis of Majorana strings. A Majorana string in canonical form is $\Gamma=\gamma_{i_{1},\sigma_{1}}\gamma_{i_{2},\sigma_{2}}\dots\gamma_{i_{l},\sigma_{l}}$, where $\gamma_{2j-1,\sigma}=c_{j\sigma}+c_{j\sigma}^{\dagger}$ and $\gamma_{2j,\sigma}=-i(c_{j\sigma}-c_{j\sigma}^{\dagger})$. The canonical ordering is fixed by (i) strictly increasing Majorana indices $i_{1}<i_{2}<\dots<i_{l}$ and (ii) for each site, the spin-down operator is placed before the spin-up operator, i.e., $\gamma_{j,\downarrow}\gamma_{j,\uparrow}$. For a system of spinless (spinful) fermions on $\cal V$ sites, there are in total ${\cal D}=4^{\cal V}$ ( ${\cal D}=4^{2{\cal V}}$) canonical Majorana strings $\Gamma_{s}$, which form an operator basis $\mathcal{G}=\{\Gamma_{s}\}_{s=1}^{{\cal D}}$ orthogonal with respect to the Frobenius inner product introduced earlier. Any fermionic operator can be uniquely decomposed in this basis. Consequently, the $n$-fold commutator can be written as
$[H(V),J]^{(n)}=\sum_{k=1}^{\mathcal{N}_{n}}a_{k}\Gamma_{k},$ (11)
In computer memory, each Majorana string $\Gamma_{k}$ is stored simply as an ordered list of indices, while the coefficients $a_{k}$ may be represented using finite-precision arithmetic, infinite-precision arithmetic, or in symbolic form. Our numerical routines operate directly in the space of Majorana strings. In practice, we first express the Hamiltonians (1,2) and the observable (3) in the Majorana basis, and then iteratively construct the commutators $H,J]^{(n)}$ until memory constraints are reached.
Using infinite-precision arithmetic, the largest operators we obtained for the two-dimensional ttVV model (1) at $k=9$ contain $\mathcal{N}^{\mathrm{2D\text{-}tV}}_{9}\simeq 7\times 10^{8}$ Majorana strings and occupy approximately $43$ GB of memory, corresponding to an average of $\sim 60$ bytes per string. The number of terms $\mathcal{N}_{n}$ grows exponentially with $n$. For example, in the three-dimensional Hubbard model we observe $\mathcal{N}^{\mathrm{3D\text{-}Hub}}_{n}\sim 10^{n}$, making each subsequent moment roughly an order of magnitude more expensive to compute than the previous one.
Although our framework allows us to keep the coefficients $a_{k}$ in symbolic form and thereby construct symbolic surrogates of the operators $H(V),J]^{(n)}$, this is not required to obtain symbolic expressions for the moments $\mu_{2n}(V)$. Since $\mu_{2n}(V)$ is a polynomial in $V$ of degree $2n$, it is sufficient to compute the moments at $n+1$ distinct values of $V$ using infinite-precision arithmetic and then reconstruct the full polynomial via interpolation. This strategy trades computational time for memory and enables access to higher-order moments in memory-limited settings.
@@
We have implemented the symbolic recursion method for strongly interacting fermions on two and three dimensional lattices, focusing on $t$-$V$ and Hubbard models. The method yields moments and Lanczos coefficients of dynamic autocorrelation functions as functions of model parameters. We have verified that the universal operator growth hypothesis (UOGH) holds for locally interacting fermionic systems. It has been shown that recursion method supplemented by UOGH describes the dynamics for all time scales up to thermalization times. This gives access to transport properties. We have computed the diffusion constant in the limit of infinite temperature and the dc conductivity at temperatures above the model bandwidths.
Our work opens the door to further developments of operator-based methods in fermionic systems. A direct extension of the present results is the computation of higher-order moments by exploiting symmetries in operator space [teng2025leveraging], in the present work we made use only of translational symmetry. Another promising direction is the application of the recursion method to quench dynamics [shirokov2025quench] and to finite-temperature calculations [angelinos2025temperature].
We have also demonstrated the effectiveness of a symbolic computational workflow. In the future, one may obtain families of symbolic expressions depending on multiple parameters, such as the chemical potential, next-nearest-neighbour hopping amplitudes, or interaction strengths, for a variety of lattice geometries. A single upfront computational investment can then yield compact symbolic representations that allow for efficient evaluation of dynamical and transport quantities, across wide regions of parameter space.

### v2

@@ Abstract
In Fig. 3, we present the infinite-temperature diffusion constants for two- and three-dimensional $t$-$V$ and Hubbard models as functions of the interaction strength $V$ (calculations are detailed in [1]). Perturbation theory predicts $D\sim 1/V^{2}$ at small $V$ [42, 90]. Surprisingly, we find that this scaling extends to quite large values of $V$, far beyond the expected range of applicability of the perturbation theory.
There exists a compelling experimental [19] and theoretical [19, 56, 88, 42, 90] evidence that the linear scaling of conductivity with $\beta$, as expressed in Eq. (10), remains valid down to relatively low temperatures on the order of 1 or even less. Combined with this observation, our calculations yield results for the dc conductivity across a wide range of interaction strengths and temperatures.
Our result for the diffusion constant in the 2D Hubbard model is consistent with a recent tour de force perturbative calculation reported in Ref. [43]. Notably, the latter study has culminated several earlier attempts [42, 90] that had missed certain relevant leading-order perturbative contributions. The intricacy of this perturbative calculation underscores the nontrivial nature of the hight-temperature quantum transport.
@@
We have implemented the symbolic recursion method for strongly interacting fermions on two and three dimensional lattices, focusing on $t$-$V$ and Hubbard models. The method yields moments and Lanczos coefficients of dynamic autocorrelation functions as symbolic computer formulas involving model parameters. We have verified that the Lanczos coefficients for the above locally interacting fermionic models are generally consistent with the universal operator growth hypothesis (UOGH). It has been demonstrated that the recursion method supplemented by UOGH is capable of computing dynamical correlation functions up to thermalization times and accurately predicting transport properties. As an illustration, we have computed the diffusion constant in the limit of infinite temperature and the dc conductivity at temperatures above the model bandwidths. Quite unexpectedly, we have found that the diffusion constant obeys a simple scaling law over a broad range of interaction strengths.
Our work paves the way for further developments of operator-based methods in fermionic systems. Immediate extensions include reducing computational complexity by exploiting additional symmetries [77] – beyond the translational symmetry employed here – as well as applying the method to quench dynamics [48, 74] and extending it to finite temperatures [3].
Our work also demonstrates the remarkable effectiveness of a symbolic computational workflow: a single upfront computational investment yields a reusable symbolic formula that allows for subsequent efficient evaluation of dynamical and transport quantities for arbitrary model parameters. This paradigm, bridging analytic insights with computational scalability, opens new pathways for systematically exploring correlated quantum matter beyond the reach of conventional numerical methods.
@@ Acknowledgments.
Data availability.
The core data required to reproduce our results – the symbolic moments (5) – is publicly available [75].
I End Matter
Here, we discuss technical aspects of computing commutators in the basis of Majorana strings and the moments (5). To compute the $n$-fold nested commutator $H(V),J]^{(n)}$, we employ an operator basis of Majorana strings. A Majorana string is a product of Majorana operators in the canonical order, $\Gamma=\gamma_{i_{1},\sigma_{1}}\gamma_{i_{2},\sigma_{2}}\dots\gamma_{i_{l},\sigma_{l}}$, where $\gamma_{2j-1,\sigma}=c_{j\sigma}+c_{j\sigma}^{\dagger}$ and $\gamma_{2j,\sigma}=-i(c_{j\sigma}-c_{j\sigma}^{\dagger})$. The canonical order implies that (i) the site labels are arranged in the strictly increasing order, $i_{1}<i_{2}<\dots<i_{l}$ and (ii) for each site, the spin-down operator is placed before the spin-up operator, i.e., $\gamma_{j,\downarrow}\gamma_{j,\uparrow}$. For a system of spinless (spinful) fermions on $\cal V$ sites, there are in total ${\cal D}=4^{\cal V}$ ( ${\cal D}=4^{2{\cal V}}$) canonical Majorana strings $\Gamma_{s}$, which form an operator basis $\mathcal{G}=\{\Gamma_{s}\}_{s=1}^{{\cal D}}$ orthogonal with respect to the inner product introduced earlier.
Any fermionic operator can be uniquely decomposed in this basis of Majorana strings. Consequently, the $n$-fold commutator can be written as
$[H(V),J]^{(n)}=\sum_{k=1}^{\mathcal{N}_{n}}a_{k}\Gamma_{k},$ (11)
In computer memory, each Majorana string $\Gamma_{k}$ is stored as an ordered list of indices, while the coefficients $a_{k}$ may be represented using finite-precision arithmetic, infinite-precision arithmetic, or in symbolic form. Our numerical routines operate directly in the space of Majorana strings. In practice, we first express the Hamiltonians (1,2) and the observable (3) in the Majorana basis, and then iteratively construct the commutators $H,J]^{(n)}$ until memory constraints are reached. Note that managing Majorana strings for fermions is strongly analogous to managing Pauli strings for qubits but requires the additional step of maintaining the canonical operator ordering, which incurs extra computational overhead. This extra step increases computational time as each string must be ordered, but does not increase the memory required.
Using infinite-precision arithmetic, the largest operators we obtained for the two-dimensional ttVV model (1) at $k=9$ contain $\mathcal{N}^{\mathrm{2D\text{-}tV}}_{9}\simeq 7\times 10^{8}$ Majorana strings and occupy approximately $43$ GB of memory, corresponding to an average of $\sim 60$ bytes per string. The number of terms $\mathcal{N}_{n}$ grows exponentially with $n$, see Fig. 4. For example, in the three-dimensional Hubbard model we observe $\mathcal{N}^{\mathrm{3D\text{-}Hub}}_{n}\sim 10^{n}$, making each subsequent moment roughly an order of magnitude more expensive to compute than the previous one.
Figure 4: Number of terms $\mathcal{N}_{n}$ in the $n$-th nested commutator (11) as a function of the commutator order $n$ for the ttVV and Hubbard models in one, two, and three dimensions.
Although our framework allows us to keep the coefficients $a_{k}$ in symbolic form and thereby construct symbolic surrogates [32, 68] of the operators $H(V),J]^{(n)}$, this is not required to obtain symbolic expressions for the moments $\mu_{2n}(V)$. Since $\mu_{2n}(V)$ is a polynomial in $V$ of degree $2n$, it is sufficient to compute the moments at $2n+1$ distinct values of $V$ using infinite-precision arithmetic and then reconstruct the full polynomial via interpolation. This strategy trades computational time for memory and enables access to higher-order moments in memory-limited settings.
Symbolic Majorana strings are also utilized in a recently proposed Majorana propagation method [54]. This method mostly focuses on simulating fermionic quantum circuits, but can also be applied to simulations of real-time dynamics by approximating the actual Heisenberg evolution by discrete Trotter steps. Recent applications of the method to small finite two-dimensional Hubbard models show that Majorana propagation is efficient in the case of small interaction strength $V$ [31], but can face limitations for larger values of $V$ [31, 2].